The following observations have been arranged in ascending order . If the median of the data is 63, find the value of "x".
29, 32, 48, 50, x, x+2, 72, 78, 84, 95

Answer by stanbon(75887)

The following observations have been arranged in ascending order . If the median of the data is 63, find the value of "x".
29, 32, 48, 50, x, x+2, 72, 78, 84, 95
------
Since there are an even number of items,
the median is the average of the middle 2
elements.
-------
(x + x+1)/2 = 63
----
2x+1 = 126
2x = 125
x = 62.5
x+1 = 63.5
================
Cheers,
Stan H.
================